OpenStudy (anonymous):

two triangles are similar and have a ratio similarity of 2:5 what is the ratio of their perimeters and the ratio for the area.

OpenStudy (lgbasallote):

the ratio of perimeters is equal to the ratio of sides the ratio of areas is equal to the SQUARE of the ratio of the sides

OpenStudy (lgbasallote):

do you get that @y0urpointismo0t ? would you like me to prove it? lol

OpenStudy (anonymous):

kind of, can you?

OpenStudy (lgbasallote):

let's say we have triangles...the side of one triangle is 2 and the side of the other is 5..for simplicity sake ratio of lengths = 2: 5 perimeter of triangle 1: 2 + 2 + 2 = 6 perimeter of triangle 2: 5 + 5 + 5 = 15 ratio of perimeters = 6 : 15 if i simplify that by dividing both by 3 i get ratio of perimeters = 2: 5 <---proven area of triangle 1: \(\sqrt 3/2\)(i wont go through details abt that im just here to prove) area of triangle 2: \(\frac{25\sqrt 3}{8}\) ratio of areas = \(\frac{\sqrt 3}{2} : \frac{25\sqrt 3}{8}\) if i rewrite it in this form: \[\large \frac{\frac{\sqrt 3}{2}}{\frac{25\sqrt 3}{8}}\] sqrt 3 and 2 will cancel out and give us \[\large \frac{1}{\frac{25}{4}}\] simplifying this..we get \(\frac{4}{25}\) in ratio form this is 4 : 25 4 is square of 2 and 25 is square of 5..thus proven @y0urpointismo0t